OpenStudy (anonymous):

Use the formula v1 x v2 = absltvalue v1 x absltvalue v2 to find the cosine of the angle between the two vectors (-7,8) and (1,0).

OpenStudy (misty1212):

hi!!

OpenStudy (misty1212):

you got the length of each of these?

OpenStudy (misty1212):

dot product should be easy since because of the \((1,0)\) it is just \(-7\)

OpenStudy (anonymous):

So Im just curious in the equation when it says multiply v1 by v2 does that mean you multiply the x and y values of (-7,8) and (1,0) together and than add the x and y values together?

OpenStudy (dumbcow):

^^ yes \[v_1*v_2 = (x_1,y_1) * (x_2,y_2) = x_1 *x_2 + y_1*y_2\]

OpenStudy (anonymous):

so the equation would come out to be -7=7costheta. Which would then simplify down to -1=costheta. Then you solve for theta giving you 180 degrees?

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

equation is \[\cos \theta = \frac{v_1 *v_2}{|v_1| |v_2|}\] \[\cos \theta = \frac{-7}{\sqrt{113}}\] \[\theta \approx 131.2\]